Grawburg Inc. maintains a call center to take orders, answer questions, and handle complaints. The costs of the call center for a number of recent months are listed below:

Calls Taken

Call Center Cost

April

7560

$

81,960

May

7529

$

81,773

June

7570

$

82,025

July

7568

$

81,997

August

7535

$

81,816

September

7549

$

81,895

October

7592

$

82,156

November

7579

$

82,092

Management believes that the cost of the call center is a mixed cost that depends on the number of calls taken.

The excel least squares regression output is as follows:

xid-11082791_1

Using the output above identify the following:

What is the “measure of goodness of fit”? (round to 2 decimal place)

What is the variable cost per unit based on the regression equation? (round to 2 decimal place)

$

What is the fixed cost portion of Inspection Costs? (round to 2 decimal place)

$

Cost equation for Inspection Costs (in the linear equation format)

Y=

Estimated Inspection costs when 973 units are produced (round to nearest whole dollar)

$
Answer #1

SOLUTION:-

Given the excel least squares regression output is as follows:-

By using regression tool in excel,we can get regression stats

1 SUMMARY OUTPUT 3 Regression Statistics 4 Multiple R 0.998197941 5 R Square 0.996399129 Adjusted R Square 0.995798984 7 Stan

1) Measure of goodness of fit = R

Square =0.996 or 1

2) Variable cost per unit = X Variable

1 = $6.09 per unit (coefficients)

3) Fixed cost portion = Intercept

= $3591.19

4) Cost equation = 6.09 * X + $35914.19

5) Estimated Inspection costs = 6.09 * X + $35914.19

= 6.09 * 973 + $35914.19

6) Estimated Inspection costs = $41839.76.

Hence proved all the Estimated inspections costs.....

THANK YOU, if any queries please leave your valuable comment on comment box......